Aspirin can be prepared from salicylic acid (C7H6O3CX7HX6OX3), which has a molar mass of 138.12 g/mol, and acetic anhydride (C4H6O3), which has a molar mass of 102.04 g/mol. The density of acetic anhydride is 1.082 g/mL.

C7H6O3+C4H6O3⟶C9H8O4+C2H3O2

What is the theoretical yield of aspirin (C9H8O4), which has a molar mass of 180.15 g/mol, possible when reacting 3.16 g of salicylic acid with 3.57 mL of acetic anhydride?
Answer #1

Density of acetic anhydride = Mass of acetic anhydride/Volume

Mass of acetic anhydride = Density * Volume = 1.082 g/ml * 3.57 ml = 3.86 grams

Number of moles of acetic anhydride = Mass/molar mass = 3.86/102.04 moles

Number of moles of C7H6O3= Mass/molar mass = 3.16/138.12 moles

Hence, C7H6O3 is the limiting reagent.

Theoretical yield of aspirin = (Number of moles) * molar mass = 3.16/138.12 * 180.15 = 4.12 grams

Hence the theoretical yield of aspirin is equal to 4.12 grams

Note - Post any doubts/queries in comments section.